Please check my answers
Please check my work
add,subtract,multiply or divide the polynomials
3x(-4x^2+5x+14) = -12x^3+15x^2+42x
(7x-5)^2 = 49x^2+25
(3x+4y)(3x-4y) = 9x^2-16y^2
(4x^2-17x-2)divided by (2x+3) =2x+23x-2/2x+3
(2/3x^3-1/2x+1/4)-(2/5x^2-1/3x+1/9) =2/3x^3-1/6x+5/26-2/5x^2
(4x^3+19x^2-7x-14) divided by(3x+4)= 4/3x^2+41/9x
(-7x^2+5xy-2y^2)+(3x^2-5xy-4y^2) = -4x^2-6y^2
(6r^3-8rs+4s^2)-(-5r^3+3rs+s^2) = 11r^2-11rs+2r^2
(8x^2+11xy-17y^2)+(2x^2+4xy-18y^2) = 10x^2+15xy-35y^2
(-4x^2y-2z^2y)-(8x^2y-4z^2y)+(-6x^2y+14xy^2) = -18x^2y=2z^2y+14xy^2
Thank you so much

Answer by checkley77(12844)

(2/3 x^3 - 1/2 x + 1/4)-(2/5 x^2 - 1/3 x + 1/9) please do that one again.
OK. I'll try.
2x^3/3-x/2+1/4-2x^2/5+x/3-1/9
2x^3/3-2x^2/5-x/2+x/3-1/9+1/4
2x^3/3-2x^2/5-(3x-2x)/2*3-(-4+9)/9*4
2x^3/3-2x^2/5-x/6-5/36
IF YOU NEED TO COMBINE THE FRACTIONS THEN THE COMMON DENOMINATOR IS 36*5=180:
(60*2x^3-36*2x^2-30*x-5*5)/180
120x^3-72x^2-30x-25)/180
------------------------------------------------------------------------
{used to show the correct values}
3x(-4x^2+5x+14) = -12x^3+15x^2+42x OK.
(7x-5)^2 = 49x^2+25 WRONG
(49x^2{-70x}+25) CORRECT.
(3x+4y)(3x-4y) = 9x^2-16y^2 OK.
(4x^2-17x-2)divided by (2x+3) =2x+23x-2/2x+3 WRONG.
(2x-{11.5) & 32.5/(2x+3)}CORRECT.
(2/3x^3-1/2x+1/4)-(2/5x^2-1/3x+1/9) =2/3x^3-1/6x+5/26-2/5x^2 WRONG.
2x^3/3-2x^2/5-x/2-x/3+1/4+1/9=2x^3/3-2x^2/5-(3x+2x)/6+(9+4)/36
2x^3/3-2x^2/5-{5x/6+13/36} CORRECT.
(4x^3+19x^2-7x-14) divided by(3x+4)= 4/3x^2+41/9x WRONG.
4x^2/3+41x/9-{227/27 & (19 17/27)/(3x+4)} CORRECT.
(-7x^2+5xy-2y^2)+(3x^2-5xy-4y^2) = -4x^2-6y^2 OK.
(6r^3-8rs+4s^2)-(-5r^3+3rs+s^2) = 11r^2-11rs+2r^2 WRONG.
11r^3-11rs+2{s^}2 CORRECT.
(8x^2+11xy-17y^2)+(2x^2+4xy-18y^2) = 10x^2+15xy-35y^2 OK.
(-4x^2y-2z^2y)-(8x^2y-4z^2y)+(-6x^2y+14xy^2) = -18x^2y=2z^2y+14xy^2 WRONG.
-18x^2y{+}2z^2y+14xy^2 CORRECT.